OpenStudy (anonymous):

One day, Bob went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

u + o = 74 2 u + 4 o = 196 - ( 2u + 2 o = 148) 0 u + 2 o = 48 2o = 48 o = 24 u = 50

OpenStudy (anonymous):

Let human, h and L be the number of humans, heads and legs respectively. Solve the following equation for human:\[\text{human}(h+2L)+(74-\text{human})(h+4L)= 74h+196L\]\[\text{human}=\{50 \text{ heads},100\text{ legs}\},\text{horses}=\{24\text{ heads},96\text{ legs}\} \]There are 50 humans and 24 horses.